OpenStudy (anonymous):

Find the length of side of an equilateral triangle with area 10.2m^2???

OpenStudy (diyadiya):

the length is already given ..what do you have to find?

OpenStudy (anonymous):

typo, excuse me.

OpenStudy (diyadiya):

Okay \[\text{Area of equilateral \triangle }= side^2 \times \frac{\sqrt{3}}{4}\]

OpenStudy (diyadiya):

Here Area is given and let the side be x now solve for x

OpenStudy (anonymous):

sorry, can we back up for a a minute.

OpenStudy (anonymous):

What happened to 1/2 b*h??

OpenStudy (diyadiya):

1/2 * b* h is the area of a Right triangle (one of the angle is \(90^o\) )

OpenStudy (anonymous):



OpenStudy (anonymous):

do I need to do something like that?

OpenStudy (diyadiya):

No ,you don't need a figure to solve this, particular, problem.

OpenStudy (diyadiya):

Here Area=10.2 \(m^2\) Formula:\[\text{Area of equilateral \triangle }= side^2 \times \frac{\sqrt{3}}{4}\] let the length of the side be \(x\) m \[\large 10.2 = x^2 \times \frac{\sqrt{3}}{4}\]

OpenStudy (diyadiya):

Understood this ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks.

OpenStudy (diyadiya):

You're Welcome :) Hope you can do the rest !!

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

can you make the calcule ?